# Policy Bifurcation in Safe Reinforcement Learning

### v2

@@ Discussion
Our study has revealed the inherent limitations of continuous policies in constrained OCPs. Specifically, our main theoretical results establish that continuous policies fail to achieve optimality and feasibility when the system’s constrained set is non-simply connected. If the initial state set is noncontractible while the target state set is contractible, no feasible continuous policy exists for the constrained OCP. Although our analysis concentrates on these topological conditions, they are not rare exceptions but rather common occurrences in a wide range of real-world control tasks. This underscores the need for alternative approaches in the design of control policies for constrained OCPs.
@@ Methods
Given the control system defined by the dynamic function $f$ and the control policy $\pi$, we define the continuous-time state transition function $F_{\pi}:\mathcal{X}\times\mathbb{R}_{\geq 0}\to\mathcal{X}$. $F_{\pi}(x_{\mathrm{init}},t)$ maps an initial state $x_{\mathrm{init}}$ to its corresponding state at a later time $t$ under the policy $\pi$. It is defined by the solution to the initial value problem:
@@
$F_{\pi}(x_{\mathrm{init}},t)=x(t)|_{x(0)=x_{\mathrm{init}}}$
@@
To facilitate subsequent proofs in the context of control systems, we consider augmenting the system’s state by incorporating time. Specifically, time $t$ is introduced as an additional dimension to the system state, leading to the augmented state denoteed as $\tilde{x}=(x,t)$. The augmented state space is defined as
@@
Consider the set of initial states $\mathrm{X}_{\mathrm{init}}$ in the context of a control system with an state transition function $F_{\pi}$. For a given finite time $T$, the reachable tuple $\mathcal{R}$ is defined as the collection of all augmented states traversed within the time interval $[0,T]$ starting from any initial state in $\mathrm{X}_{\mathrm{init}}$. Formally,
@@
In this work, we assume that the state space of the constrained OCP is a subset of a Euclidean space. Since Euclidean spaces are topological spaces with the standard topology induced by the Euclidean metric, any subset of a Euclidean space, including our state space, can be considered a topological space. Therefore, when we refer to a set of state space, it is understood to be equipped with the topology inherited from the Euclidean space.
@@
The concept of path and loop is important for the concept of ‘non-simply connected’ and ‘contractibility of a subset’. ‘non-simply connected’ is a fundamental concept in topology, referring to a connected space that contains loops which cannot be continuously contracted to a point, usually due to obstacles. The ‘contractibility of a subset’ is further detailed in Definition 5. These concepts, illustrated in Fig. 6, are essential for analyzing the topological properties of the reachable tuple and assessing the feasibility of continuous policies.
Figure 6: Visualization of topological concepts. This figure illustrates the concepts of contractibility as defined in Definition 5. The space $\mathrm{X}_{\mathrm{cstr}}$ is non-simply connected due to the presence of $\mathrm{X}_{\mathrm{viol}}$. The initial state space $\mathrm{X}_{\mathrm{init}}$ is noncontractible within $\mathrm{X}_{\mathrm{cstr}}$, as shown by the loop $\ell_{\mathrm{init}}$, which cannot be continuously contracted to a point within $\mathrm{X}_{\mathrm{cstr}}$. In contrast, the target state space $\mathrm{X}_{\mathrm{target}}$ is an obstacle-free region, allowing every loop to be continuously contracted to a point, demonstrating its contractibility within $\mathrm{X}_{\mathrm{cstr}}$.
@@
    Finite Time Reachability: $F_{\pi}(x(0),T)\in\mathrm{X}_{\mathrm{target}}$ for all $x(0)\in\mathrm{X}_{\mathrm{init}}$.
@@
    Target Set Contractibility: $\mathrm{X}_{\mathrm{target}}$ is contractible.
@@
Given the Finite Time Reachability (condition 4) and the fact that $\ell_{\mathcal{R}}(s)\in\mathcal{R}$, it follows that $F_{\pi}(\phi_{x}(\ell_{\mathcal{R}}(s)),T-\phi_{t}(\ell_{\mathcal{R}}(s)))\in\mathrm{X}_{\mathrm{target}}$. Consequently, $\ell_{\mathcal{R},T}^{x}$ forms a loop in $\mathrm{X}_{\mathrm{target}}$.
According to the Target Set Contractibility (condition 5), a continuous mapping $H_{c}:[0,1]\times[0,T]\to\mathrm{X}_{\mathrm{cstr}}$ exists such that $H_{c}(s,0)=\ell_{\mathcal{R},T}^{x}(s)$, and there exists a constant $c\in\mathrm{X}_{\mathrm{target}}$ such that $H_{c}(s,T)=c$ for all $s\in[0,1]$.
@@
If the optimal solution of an open-loop constrained OCP corresponds to a reachable tuple $\mathcal{R}$ that cannot be contracted within $\tilde{\mathrm{X}}_{\mathrm{cstr}}$, this optimal solution will be discontinuous. Consequently, continuous policies fail to attain this optimal solution of the open-loop constrained OCP.
@@
We will prove the theorem by contradiction. According to Lemma 2, if a continuous policy $\pi_{\mathrm{conti}}$ is feasible, then its corresponding reachable tuple $\mathcal{R}$ is contractible. This implies that if the reachable tuple $\mathcal{R}$ of $\pi_{\mathrm{conti}}$ is noncontractible, then $\pi_{\mathrm{conti}}$ is not feasible. Therefore, for an open-loop OCP whose optimal solution leads to a reachable tuple $\mathcal{R}$ that is noncontractible, the optimal solution cannot be represented in the form of a continuous policy function. Consequently, continuous policies fail to attain this optimal solution of the open-loop constrained OCP. ∎
@@
The desired continuous policy is denoted as $\pi_{\mathrm{des}}$. Under $\pi_{\mathrm{des}}$, the vehicle is expected to: (1) circle right from $\tilde{x}_{\mathrm{a}}$ to reach $\tilde{x}_{\mathrm{a^{\prime}}}$ and circle left from $\tilde{x}_{\mathrm{b}}$ to reach $\tilde{x}_{\mathrm{b^{\prime}}}$; (2) ensure that, starting from any state $x\in\mathrm{X}_{\mathrm{init}}$, it avoids collisions with the obstacle and reaches the target set within finite time. We will demonstrate, using Lemma 2, that these two conditions present a contradiction for a continuous policy.
To formalize the trajectories under the desired policy $\pi_{\mathrm{des}}$, we define paths $\ell_{\mathrm{aa^{\prime}}}(s)$ and $\ell_{\mathrm{bb^{\prime}}}(s)$. These paths represent the trajectories of the vehicle under the policy $\pi_{\mathrm{des}}$ from the initial states $\tilde{x}_{\mathrm{a}}$ and $\tilde{x}_{\mathrm{b}}$ to their respective target states $\tilde{x}_{\mathrm{a^{\prime}}}$ and $\tilde{x}_{\mathrm{b^{\prime}}}$ over a normalized time scale $s\in[0,1]$. Specifically, the paths are defined as
@@
$\ell_{\mathrm{ab}}$ is a path in $\mathrm{X}_{\mathrm{init}}$, it starts from $\tilde{x}_{\mathrm{a}}$, ends at $\tilde{x}_{\mathrm{b}}$. $\ell_{\mathrm{a^{\prime}b^{\prime}}}$ is a path in $\mathrm{X}_{\mathrm{target}}$, it starts from $\tilde{x}_{\mathrm{a^{\prime}}}$, ends at $\tilde{x}_{\mathrm{b^{\prime}}}$. Consider the loop $\ell_{aa^{\prime}b^{\prime}ba}$ within $\mathcal{R}$, which transitions through the sequence $a\to a^{\prime}\to b^{\prime}\to b\to a$:
@@
Consider a constrained OCP characterized by a Lipschitz continuous dynamic function $f$ and policy $\pi$. The problem has no feasible solution if the initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible, and the target state set $\mathrm{X}_{\mathrm{target}}$ is contractible. Specifically, this constrained OCP is unsolvable when the following six conditions are simultaneously met:
@@
    Finite Time Reachability: $F_{\pi}(x(0),T)\in\mathrm{X}_{\mathrm{target}}$ for all $x(0)\in\mathrm{X}_{\mathrm{init}}$.
@@
    Target Set Contractibility: $\mathrm{X}_{\mathrm{target}}$ is contractible.
@@
As established in Lemma 2, under conditions 1 to 5, the reachable tuple $\mathcal{R}$ is contractible. However, we will now demonstrate, using Initial State Set Noncontractibility (condition 6), that this contractibility does not hold for $\mathcal{R}$ within $\mathrm{\tilde{X}}_{\mathrm{cstr}}$.
According to the definition of contractibility (see Definition 5), condition 6 implies the existence of a loop $\ell_{\mathrm{init}}:[0,1]\to\mathrm{X}_{\mathrm{init}}$ such that there is no continuous mapping $H:[0,1]\times[0,T]\to\mathrm{X}_{\mathrm{cstr}}$ for which $H(s,0)=\ell_{\mathrm{init}}(s)$ and $H(s,T)=c$ for some $c\in\mathrm{X}_{\mathrm{cstr}}$ and all $s\in[0,1]$. This indicates that $\ell_{\mathrm{init}}$ cannot be continuously contracted to a point within $\mathrm{X}_{\mathrm{cstr}}$.
The augmented loop $\ell_{\mathrm{init},\mathrm{aug}}(s)=(\ell_{\mathrm{init}}(s),0)$ forms a loop in $\mathcal{R}$. To demonstrate the non-existence of a continuous mapping capable of contracting $\ell_{\mathrm{init},\mathrm{aug}}$ to a point within $\mathrm{\tilde{X}}_{\mathrm{cstr}}$, let’s assume, for the sake of contradiction, that a continuous mapping $H_{\mathrm{init},\mathrm{aug}}$ exists. Then, define $H_{\mathrm{init}}:[0,1]\times[0,T]\to\mathrm{X}_{\mathrm{cstr}}$ as $H_{\mathrm{init}}(s,t)=\phi_{x}(H_{\mathrm{init},\mathrm{aug}}(s,t))$. $H_{\mathrm{init}}$ contracts the loop $\ell_{\mathrm{init}}$ to a point in $\mathrm{X}_{\mathrm{cstr}}$. This, however, directly conflicts with the noncontractibility of $\mathrm{X}_{\mathrm{init}}$ as stated in condition 6. Hence, this contradiction demonstrates that $\mathcal{R}$ is noncontractible under condition 6, since $\ell_{\mathrm{init},\mathrm{aug}}$ cannot be contracted to a point within $\mathrm{\tilde{X}}_{\mathrm{cstr}}$.
Summarizing, conditions 1 to 5 imply that the reachable tuple $\mathcal{R}$ is contractible. However, condition 6 leads to the noncontractibility of $\mathcal{R}$. This contradiction underlines the insolvability of the constrained OCP under these specific conditions. ∎
In summary, continuous policies fail to achieve optimality and feasibility in scenarios with complex constraints, indicating the need for an improved policy structure. The following sections explore Safe RL with bifurcated policies to address the challenges inherent in continuous policies.
@@
Through this method, the stochastic policy with Gaussian mixture distribution is transformed into a bifurcated policy, characterized by discontinuous action outputs that change abruptly depending on the state $x$. The training process, discussed in subsequent sections, involves learning the parameters $\theta$ to effectively implement this policy in varying environments.
@@
In this section, we propose the Multimodal Policy Optimization (MUPO), a safe RL algorithm specifically designed for the efficient learning of bifurcated policy. Built upon the Actor-Critic architecture, MUPO comprises an actor learning a bifurcated policy $\pi_{\theta}$ and a critic evaluating the policy. The critic utilizes two distinct action-value distribution functions denoted as $\mathcal{Z}_{\omega_{1}}$ and $\mathcal{Z}_{\omega_{2}}$, where $\omega_{1}$ and $\omega_{2}$ denote their parameters respectively.
@@
Our algorithm incorporates action-value distribution from the DSAC algorithm [22] within the critic component, enabling a comprehensive capture of potential future rewards. This integration is particularly crucial given the multimodal nature of our policy function, which often results in a significant variance in state-action values (i.e., Q-values). Compared to methods like soft actor-critic (SAC) [19] and proximal policy optimization (PPO) [10], which utilize deterministic value functions, the action-value distribution in DSAC algorithm effectively mitigate the overestimation issue in value functions, thereby enhancing policy performance. The loss function for the action-value distribution $\mathcal{Z}_{\omega}$ is defined as the negative expected log-likelihood of the target value given the current value distribution:
@@
Policy optimization aims at aligning the current policy $\pi_{\theta}$ as closely as possible with the energy-based policy $\pi_{\omega}$. While various projection methods are feasible, utilizing the Kullback-Leibler (KL) divergence proves to be particularly effective.
The KL divergence is a measure of the difference between two probability distributions, it comes in two forms: forward and reverse. The reverse KL divergence, defined as $\mathbb{E}_{x\sim\mathcal{B}}\left[D_{\mathrm{KL}}(\pi_{\theta}(\cdot|x)\|\pi_{\omega}(\cdot|x))\right]$, is highly efficient for uni-modal policies such as Gaussian distributions. However, in multi-modal distribution scenarios, reverse KL divergence tends to learn only one mode while neglecting others, limiting its effectiveness in capturing the full complexity of such distributions.
On the other hand, the forward KL divergence, expressed as $\mathbb{E}_{x\sim\mathcal{B}}\left[D_{\mathrm{KL}}(\pi_{\omega}(\cdot|x)\|\pi_{\theta}(\cdot|x))\right]$, plays an important role in learning multi-modal distributions. It fosters a wider coverage by the policy, which is essential for exploring diverse behavioral modes [29]. However, the accurate computation of forward KL divergence relies on precise sampling from the energy-based policy $\pi_{\omega}$, which is computationally challenging. Consequently, policies learned using forward KL divergence may not achieve high performance.

### v3

@@ Discussion
Our study has revealed the inherent limitations of continuous policies in constrained OCPs. Specifically, our main theoretical results establish that continuous policies fail to achieve optimality and feasibility when the system’s constrained set is non-simply connected. If the initial state set is noncontractible while the goal state set is contractible, no feasible continuous policy exists for the constrained OCP. Although our analysis concentrates on these topological conditions, they are not rare exceptions but rather common occurrences in a wide range of real-world control tasks. This underscores the need for alternative approaches in the design of control policies for constrained OCPs.
@@ Methods
Given the control system defined by the dynamic function $f$ and the control policy $\pi$, we define the continuous-time state transition function as $F_{\pi}:\mathcal{X}\times\mathbb{R}_{\geq 0}\to\mathcal{X}$. $F_{\pi}(x_{\mathrm{init}},t)$, which maps an initial state $x_{\mathrm{init}}$ to the subsequent state at a later time $t$ under the policy $\pi$. It is determined by solving the initial value problem:
@@
$F_{\pi}(x_{\mathrm{init}},t)=x(t)|_{x(0)=x_{\mathrm{init}}}.$
@@
To facilitate subsequent proofs in the context of control systems, we consider augmenting the system’s state by incorporating time. Specifically, time $t$ is introduced as an additional dimension to the system state, leading to the augmented state denoted as $\tilde{x}=(x,t)$. The augmented state space is defined as
@@
Consider the set of initial states $\mathrm{X}_{\mathrm{init}}$ in the context of a control system with a state transition function $F_{\pi}$. For a given finite time $T$, the reachable tuple $\mathcal{R}$ is defined as the collection of all augmented states traversed within the time interval $[0,T]$ starting from any initial state in $\mathrm{X}_{\mathrm{init}}$, i.e.,
@@
In this work, we assume that the state space of the constrained OCP is a subset of Euclidean space. Since Euclidean spaces are topological spaces with the standard topology induced by the Euclidean metric, any subset of Euclidean space, including our state space, can be considered a topological space. Therefore, when we refer to a set of the state space, it is understood to be equipped with the topology inherited from the Euclidean space.
@@
The concept of path and loop is important for the concept of “non-simply connected” and “contractibility of a subset”. “non-simply connected” is a fundamental concept in topology, referring to a connected space that contains loops that cannot be continuously contracted to a point, usually due to obstacles. The “contractibility of a subset” is further detailed in Definition 5. These concepts, illustrated in Fig. 6, are essential for analyzing the topological properties of the reachable tuple and assessing the feasibility of continuous policies.
Figure 6: Visualization of topological concepts. This figure illustrates the concepts of contractibility as defined in Definition 5. The space $\mathrm{X}_{\mathrm{cstr}}$ is non-simply connected due to the presence of $\mathrm{X}_{\mathrm{viol}}$. The initial state space $\mathrm{X}_{\mathrm{init}}$ is noncontractible within $\mathrm{X}_{\mathrm{cstr}}$, as shown by the loop $\ell_{\mathrm{init}}$, which cannot be continuously contracted to a point within $\mathrm{X}_{\mathrm{cstr}}$. In contrast, the goal state space $\mathrm{X}_{\mathrm{goal}}$ is an obstacle-free region, allowing every loop to be continuously contracted to a point, demonstrating its contractibility within $\mathrm{X}_{\mathrm{cstr}}$.
@@
    Finite time reachability: $F_{\pi}(x(0),T)\in\mathrm{X}_{\mathrm{goal}}$ for all $x(0)\in\mathrm{X}_{\mathrm{init}}$.
@@
    Goal set contractibility: $\mathrm{X}_{\mathrm{goal}}$ is contractible.
@@
Given the finite time reachability (condition 4) and the fact that $\ell_{\mathcal{R}}(s)\in\mathcal{R}$, it follows that $F_{\pi}(\phi_{x}(\ell_{\mathcal{R}}(s)),T-\phi_{t}(\ell_{\mathcal{R}}(s)))\in\mathrm{X}_{\mathrm{goal}}$. Consequently, $\ell_{\mathcal{R},T}^{x}$ forms a loop in $\mathrm{X}_{\mathrm{goal}}$.
According to the goal set contractibility (condition 5), a continuous mapping $H_{c}:[0,1]\times[0,T]\to\mathrm{X}_{\mathrm{cstr}}$ exists such that $H_{c}(s,0)=\ell_{\mathcal{R},T}^{x}(s)$, and there exists a constant $c\in\mathrm{X}_{\mathrm{goal}}$ such that $H_{c}(s,T)=c$ for all $s\in[0,1]$.
@@
If the optimal solution of an open-loop constrained OCP corresponds to a reachable tuple $\mathcal{R}$ that cannot be contracted within $\tilde{\mathrm{X}}_{\mathrm{cstr}}$, then this optimal solution will be discontinuous. Consequently, continuous policies fail to represent this optimal solution of the open-loop constrained OCP.
@@
We prove this theorem by contradiction. According to Lemma 2, if a continuous policy $\pi_{\mathrm{conti}}$ is feasible, then its corresponding reachable tuple $\mathcal{R}$ is contractible. This implies that if the reachable tuple $\mathcal{R}$ of $\pi_{\mathrm{conti}}$ is noncontractible, then $\pi_{\mathrm{conti}}$ is not feasible. Therefore, for an open-loop OCP whose optimal solution leads to a reachable tuple $\mathcal{R}$ that is noncontractible, the optimal solution cannot be represented in the form of a continuous policy function. Consequently, continuous policies fail to represent this optimal solution of the open-loop constrained OCP. ∎
@@
The desired continuous policy, denoted as $\pi_{\mathrm{des}}$, mandates the vehicle to (1) execute a right detour from $\tilde{x}_{\mathrm{a}}$ to reach $\tilde{x}_{\mathrm{a^{\prime}}}$ and a left detour from $\tilde{x}_{\mathrm{b}}$ to reach $\tilde{x}_{\mathrm{b^{\prime}}}$; (2) ensure collision avoidance with the obstacle and achieve goal set within finite time irrespective of initial state $x\in\mathrm{X}_{\mathrm{init}}$. Using Lemma 2, we will demonstrate that these two conditions present a contradiction for a continuous policy.
To formalize the trajectories under the desired policy $\pi_{\mathrm{des}}$, we define paths $\ell_{\mathrm{aa^{\prime}}}(s)$ and $\ell_{\mathrm{bb^{\prime}}}(s)$. These paths represent the trajectories of the vehicle under the policy $\pi_{\mathrm{des}}$ from the initial states $\tilde{x}_{\mathrm{a}}$ and $\tilde{x}_{\mathrm{b}}$ to their respective goal states $\tilde{x}_{\mathrm{a^{\prime}}}$ and $\tilde{x}_{\mathrm{b^{\prime}}}$ over a normalized time scale $s\in[0,1]$. Specifically, the paths are defined as
@@
$\ell_{\mathrm{ab}}$ is a path in $\mathrm{X}_{\mathrm{init}}$, it starts from $\tilde{x}_{\mathrm{a}}$, ends at $\tilde{x}_{\mathrm{b}}$. $\ell_{\mathrm{a^{\prime}b^{\prime}}}$ is a path in $\mathrm{X}_{\mathrm{goal}}$, it starts from $\tilde{x}_{\mathrm{a^{\prime}}}$, ends at $\tilde{x}_{\mathrm{b^{\prime}}}$. Consider the loop $\ell_{aa^{\prime}b^{\prime}ba}$ within $\mathcal{R}$, which transitions through the sequence $a\to a^{\prime}\to b^{\prime}\to b\to a$:
@@
Consider a constrained OCP characterized by a Lipschitz continuous dynamic function $f$ and policy $\pi$. The problem has no feasible solution if the initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible, and the goal state set $\mathrm{X}_{\mathrm{goal}}$ is contractible. Specifically, this constrained OCP is unsolvable when the following six conditions are simultaneously met:
@@
    Finite time reachability: $F_{\pi}(x(0),T)\in\mathrm{X}_{\mathrm{goal}}$ for all $x(0)\in\mathrm{X}_{\mathrm{init}}$.
@@
    Goal set contractibility: $\mathrm{X}_{\mathrm{goal}}$ is contractible.
@@
As established in Lemma 2, under conditions 1 to 5, the reachable tuple $\mathcal{R}$ is contractible. However, by incorporating the initial state set noncontractibility (condition 6), we demonstrate that this contractibility does not hold for $\mathcal{R}$ within $\mathrm{\tilde{X}}_{\mathrm{cstr}}$.
According to the definition of contractibility (see Definition 5), condition 6 implies the existence of a loop $\ell_{\mathrm{init}}:[0,1]\to\mathrm{X}_{\mathrm{init}}$ such that there is no such continuous mapping $H:[0,1]\times[0,T]\to\mathrm{X}_{\mathrm{cstr}}$, where $H(s,0)=\ell_{\mathrm{init}}(s)$ and $H(s,T)=c$ for some $c\in\mathrm{X}_{\mathrm{cstr}}$ and all $s\in[0,1]$. This indicates that $\ell_{\mathrm{init}}$ cannot be continuously contracted to a point within $\mathrm{X}_{\mathrm{cstr}}$.
The augmented loop $\ell_{\mathrm{init},\mathrm{aug}}(s)=(\ell_{\mathrm{init}}(s),0)$ forms a loop in $\mathcal{R}$. To demonstrate the non-existence of a continuous mapping capable of contracting $\ell_{\mathrm{init},\mathrm{aug}}$ to a point within $\mathrm{\tilde{X}}_{\mathrm{cstr}}$, we initially assume that a continuous mapping $H_{\mathrm{init},\mathrm{aug}}$ exists. Then, define $H_{\mathrm{init}}:[0,1]\times[0,T]\to\mathrm{X}_{\mathrm{cstr}}$ as $H_{\mathrm{init}}(s,t)=\phi_{x}(H_{\mathrm{init},\mathrm{aug}}(s,t))$. $H_{\mathrm{init}}$ contracts the loop $\ell_{\mathrm{init}}$ to a point in $\mathrm{X}_{\mathrm{cstr}}$. However, this directly conflicts with the noncontractibility of $\mathrm{X}_{\mathrm{init}}$ as stated in condition 6. Hence, this contradiction demonstrates that $\mathcal{R}$ is noncontractible under condition 6, since $\ell_{\mathrm{init},\mathrm{aug}}$ cannot be contracted to a point within $\mathrm{\tilde{X}}_{\mathrm{cstr}}$.
In summary, conditions 1 to 5 imply that the reachable tuple $\mathcal{R}$ is contractible. However, condition 6 leads to the noncontractibility of $\mathcal{R}$. This contradiction underlines the unsolvability of the constrained OCP under these specific conditions. ∎
Consequently, continuous policies fail to achieve optimality and feasibility in scenarios with complex constraints, indicating the need for an improved policy structure. The following sections explore Safe RL with bifurcated policies to address the challenges inherent in continuous policies.
@@
Through this approach, the stochastic policy with Gaussian mixture distribution is transformed into a bifurcated policy, characterized by discontinuous action outputs that change abruptly depending on the state $x$. The training process, discussed in subsequent sections, involves learning the parameters $\theta$ to implement this policy in varying environments effectively.
@@
In this section, we propose MUPO, a safe RL algorithm specifically designed for the efficient learning of bifurcated policy. Built upon the actor-critic architecture, MUPO comprises an actor learning a bifurcated policy $\pi_{\theta}$ and a critic evaluating the policy. The critic utilizes two distinct action-value distribution functions denoted as $\mathcal{Z}_{\omega_{1}}$ and $\mathcal{Z}_{\omega_{2}}$, where $\omega_{1}$ and $\omega_{2}$ represent their parameters respectively.
@@
Our algorithm incorporates action-value distribution from the DSAC algorithm [22] within the critic component, enabling a comprehensive capture of potential future rewards. This integration is particularly crucial given the multimodal nature of our policy function, which often results in a significant variance in state-action values (i.e., Q-values). Compared to methods like SAC [19] and PPO [10], which utilize deterministic value functions, the action-value distribution in the DSAC algorithm effectively mitigates the overestimation issue in value functions, thereby enhancing policy performance. The loss function for the action-value distribution $\mathcal{Z}_{\omega}$ is defined as the negative expected log-likelihood of the target value given the current value distribution:
@@
Policy optimization aims at aligning the current policy $\pi_{\theta}$ as closely as possible with the energy-based policy $\pi_{\omega}$. While various projection methods are feasible, utilizing the KL divergence proves to be particularly effective.
The KL divergence is a measure of the difference between two probability distributions, it comes in two forms: forward and reverse. The reverse KL divergence, defined as $\mathbb{E}_{x\sim\mathcal{B}}\left[D_{\mathrm{KL}}(\pi_{\theta}(\cdot|x)\|\pi_{\omega}(\cdot|x))\right]$, is highly efficient for uni-modal policies such as Gaussian distributions. However, in multi-modal distribution scenarios, reverse KL divergence tends to learn only one mode while neglecting others [29], limiting its effectiveness in capturing the full complexity of such distributions.
On the other hand, the forward KL divergence, expressed as $\mathbb{E}_{x\sim\mathcal{B}}\left[D_{\mathrm{KL}}(\pi_{\omega}(\cdot|x)\|\pi_{\theta}(\cdot|x))\right]$, plays an important role in learning multi-modal distributions. It fosters a wider coverage by the policy, which is essential for exploring diverse behavioral modes. However, the accurate computation of forward KL divergence relies on precise sampling from the energy-based policy $\pi_{\omega}$, which is computationally challenging. Consequently, policies learned using forward KL divergence may not achieve high performance.